Plummer corporation acquired 90 percent of Softek Technologies' voting stock by issuing 200,000 shares of $1 par common stock with a fair value of $100,000,000. In addition, Plummer paid $2,000,000 in cash to the consultants and accountants who advised in the acquisition. Softek's shareholders' equity at eh date of the acquisition is as follows:

Common Stock $400,000
Additional paid-in capital $20,000,000
Retained deficit -$10,000,000
Accumulated other comprehensive loss -$1,000,000
Treasury stock -$500,000
Total $8,900,000

Softek's assets and liabilities were carried at fair value except as noted below:

Book Value Fair Value
Plant assets, net $12,000,000 $6,000,000
Trademarks $-   $2,000,000
Customer lists $-   $3,000,000

The fair value of the noncontrolling interest is estimated to be $9,000,000 at the date of acquisition.

Required:

Prepare the acquisition entry on Plummer's books and the working paper consolidation eliminating entries at the date of acquisition, following US GAAP.
Answer #1

Plummer’s acquisition entry:

Entry

Account titles and explanation

Debit

Credit

Investment in Softek

100,000,000

Merger expenses

2,000,000

Cash

2,000,000

Common stock, par value

200,000

Additional paid-in capital (100000000-200000)

99,800,000

Consolidation eliminating entries:

Entry

Account titles and explanation

Debit

Credit

Entry E

Common stock

400,000

Additional paid-in capital

20,000,000

Retained earnings

10,000,000

Accumulated OCI

1,000,000

Treasury stock

500,000

Investment in Softek (8900000*90%)

8010000

Noncontrolling interest in Softek

890000

Entry R

Trademarks

2,000,000

Customer lists

3,000,000

Goodwill

101100000

Plant assets, net (6000000-12000000)

6,000,000

Investment in Softek

91990000

Noncontrolling interest in Softek (balancing figure)

8110000

Goodwill = (100,000,000+9,000,000)-(8900000-6,000,000+2,000,000+3,000,000) = 109,000,000-7900000 = 101100000

Investment in Softek =(90%*(2000000+3000000-6000000))+(100000000-(90%*101100000)) = 91990000

Dear Student,

If any doubts, then feel free to ask to clear all your confusions.